OpenStudy (anonymous):

Simplfy, ln 4 + ln x − 1 − ln y

OpenStudy (anonymous):

\[\ln (4x/y) -1\]

OpenStudy (anonymous):

\[\ln(4)+\ln(x-1)-\ln(y)\]?